# Kareem Amin  Alex Bie  Weiwei Kong  Alexey Kurakin Natalia Ponomareva  Umar Syed  Andreas Terzis  Sergei Vassilvitskii Google

### v1

@@ 4 Privacy analysis
In this section we prove that Algorithm 1 preserves the privacy of the sensitive prompts it uses to generate synthetic examples.
@@
The proof is in Appendix C and makes use of sharp privacy analyses of: (1) zCDP to approximate DP conversion [Canonne et al., 2020]; and (2) zCDP bounds for the exponential mechanism [Cesar and Rogers, 2021].
@@ 5 Experiments
Gemma 1.1 2B IT [Team, 2024] is the data generator in all private prediction experiments. We choose it due to its lightweight, open-source JAX implementation that makes easy to implement and share sampling algorithms. [^2] https://github.com/google-deepmind/gemma Tables 1(a) and 1(b) give an overview of datasets and models used.
@@
Gemma 2B 1.1 IT Generation; private prediction
@@
(b) Overview of models used.
Table 1: Overview of datasets and models used. Datasets are benchmark classification and extraction tasks used in prior work on private synthetic text generation, with the exception of WikiMoviesJSON, which is used for structured data experiments. LaMDA and Gemma are used for synthetic data generation, while the other models are used to measure how useful our synthetic data is for improving accuracy on real test data.
@@
  • In-context learning (§5.1); We generate examples to use as in-context exemplars for prompting an LLM. We report downstream accuracy on real test examples, when prompted with synthetic data, on 3 classification tasks (AGNews [Zhang et al., 2015], DBPedia [Zhang et al., 2015], TREC [Voorhees and Tice, 2000]) and 2 extraction tasks (MIT-G, MIT-D [Liu et al., 2012]).
  • Fine-tuning (§5.2); We generate synthetic examples to use for fine-tuning a BERT classifier. We report downstream accuracy on real test examples for 3 classification tasks (IMDB [Maas et al., 2011], Yelp [Zhang et al., 2015], AGNews [Zhang et al., 2015]).
  • Structured data (§5.3); We generate examples that must adhere to structural constraints to be useful synthetic data. We consider a JSON generation task (WikiMoviesJSON [Rust, 2024]), evaluating structure preservation.
@@ 5.1 In-context learning
Using our method, we generate 90-1500 examples using Gemma 2B 1.1 IT. We compare against real examples, and results reported in the prior work of Tang et al. [2024], where they generated 4-shot examples for in-context learning. [^3] It is no longer possible to reproduce their results, due to changes in the OpenAI API since publication: GPT-3 babbage is now deprecated, and it is no longer possible to query for top 100 logprobs, which is required by their method. To evaluate generated synthetic data, we put synthetic examples in the context window before querying with the real test example, as shown in Figure 2.
1Classify the following examples: 2Text: lorem ipsum # synthetic text 1 3Answer: label 4... 5Text: sed do eiusmod # synthetic text n 6Answer: label 7 8Text: excepteur sint # test text 9Answer:
@@
We perform this evaluation with GPT-3 babbage-002 which has a 16K context window. We report results on AGNews, DBPedia, TREC, MIT-G, and MIT-D using the implementation of Zhao et al. [2021]. [^4] https://github.com/tonyzhaozh/few-shot-learning. Following the work of Tang et al. [2024], we enable contextual calibration [Zhao et al., 2021] for classification but not extraction tasks. Our evaluation setup is a best-effort reproduction of their setup, which is no longer possible to completely reproduce due to changes to OpenAI API access (see Table 2 caption for more details). Due to cost, we follow prior work [Bertsch et al., 2024, Ratner et al., 2023, Lu et al., 2022, Zhao et al., 2021] and opt to subsample test sets to 250 test examples. We run 3 seeds of label-stratified sampling of exemplars from synthetic/real data.
@@
Tang et al. [2024] 4 Tang et al. [2024]* GPT-3 babbage $69.3_{4.8}$ $82.3_{3.7}$ $50.6_{6.9}$ $54.4_{7.0}$
@@
$1$ Tang et al. [2024] 4 Tang et al. [2024]* GPT-3 babbage $64.1_{3.9}$ $81.2_{3.0}$ $50.7_{4.1}$ $46.3_{7.8}$ $69.2_{7.9}$
@@
Table 2: In-context learning results with GPT-3 babbage-002. We report mean and standard deviation over 3 random samplings (equally many from each label for classification; fully random for extraction) of synthetic/real data. (*) Note: For the results reported in Tang et al. [2024], they use GPT-3 babbage (now deprecated; we use GPT-3 babbage-002) as the in-context learner, and use the top 100 logprobs for contextual calibration (only top 5 are available now); while not directly comparable, we report their results for context.
@@ Results.
Results are presented in Table 2. Our gains in quantity while maintaining quality are realized in terms of 64-shot in-context learning accuracy. In some cases, we can generate more examples, but we limit ourselves to 64 for these evaluations for cost and efficiency reasons. Our results at 64 shots are comparable to real data at 64 shots. Notably, our synthetic data at 64 shots improves over real data at 4 shots – which is roughly an upper bound on the performance of methods limited to generating 4 examples (e.g., Tang et al. [2024]). We also improve over results reported in Tang et al. [2024], although we note that there are differences in the experimental setup.
@@ 5.2 Fine-tuning
We use our approach to generate a large quantity of synthetic data for the purposes of fine-tuning 110M BERT-Base models. We consider 3 classification tasks used in prior work on private fine-tuning [Kurakin et al., 2024]), following the exact same evaluation procedure. We omit comparison to prior private prediction work (e.g. [Tang et al., 2024]), as they only generate 4 examples which is insufficient for fine-tuning.
@@ Results.
Main results are presented in Table 3. Across various datasets and privacy levels, we generate between 2.5K (IMDB, $\varepsilon$=1) and 200K (Yelp, $\varepsilon$=10) examples for fine-tuning. Prior work generating fewer than $10$ examples using private prediction were unable to compare with private fine-tuning on these tasks at all. While there remains a gap between the best fine-tuning and best private inference methods on downstream classification tasks, we achieve reasonable performance, even out-performing or matching the baseline of privately tuning all the parameters in the model reported in Kurakin et al. [2024].
@@
Real data [Kurakin et al., 2024] - $93.7_{0.1}$ $97.6_{0.1}$ $93.7_{0.1}$
Fine-tune [Kurakin et al., 2024] LaMDA 8B $93.2_{0.2}$ $79.1_{1.7}$ $83.9_{0.6}$ $84.0_{0.7}$ $95.9_{0.1}$ $84.1_{0.3}$ $84.6_{0.1}$ $84.2_{0.3}$ $91.1_{0.1}$ $65.7_{2.9}$ $65.3_{2.7}$ $65.1_{5.3}$
@@
Table 3: Results of fine-tuning on real and synthetic data with BERT. We report mean and standard deviation over 3 runs of downstream fine-tuning and evaluation. We compare to results reported in [Kurakin et al., 2024] that fine-tunes a synthetic data generator with DP-SGD. We generate 2.5-200K examples with private prediction, which suffices to train reasonably performing models on.
@@
We additionally consider the limited data regime. In Appendix A we present experiments on AGNews1K, a 1024-subsample of AGNews. Our method, which employs parallel composition, is “pay-as-you-go”, i.e., we can put in a small amount of data to get out a small amount, while preserving quality. On the other hand, fine-tuning based approaches necessarily pay upfront to ensure the model and all future generations are private. This means that without sufficient data, all outputs will be low quality. Results in Table 5 demonstrate that our private prediction method generates more useful examples for in-context learning in this regime.
@@ 5.3 Structured data
For JSON generation, we evaluate on a dataset of information about American movies scraped from Wikipedia [Rust, 2024]. Entries contain fields such as title, year, cast, and extract (a short synopsis). We lightly curate the data: we omit uninteresting fields (i.e., thumbnail dimensions and URLs) and remove entries with incomplete entries. We refer to the resulting 34,266 JSON examples with 6 fields as WikiMoviesJSON. We evaluate two criteria: the rate at which output generated constitutes well-formed JSON (Parses (%)), and rate at which the output passes basic schema validation (Validates (%)). This includes checks such as: no extra fields, all required fields are present, values are the correct type, and other custom constraints (e.g. no whitespace in the href field).
@@ 6 Discussion
We believe that our significantly improved performance relative to Tang et al. [2024] is primarily attributable to two algorithmic innovations.
First, for each generated token, Tang et al. [2024] preserve the privacy of the entire distributions from which the token is sampled (by taking argmax), even though only the token itself is included in the synthetic data. By contrast, our method uses a discrete choosing mechanism, the exponential mechanism. As a result, we do not need to maintain a DP version of the entire token distribution to release a single token. This decision leads to significantly lower noise requirements, as a straightforward calculation reveals. Empirically, we obtained good synthetic data quality with $s=250$, $\tau=2$, $c=10$ and $\delta=10^{-6}$. In order to switch to the Gaussian mechanism using its standard $(\varepsilon,\delta)$-DP guarantee, and achieve comparable privacy guarantees we would would require $\sigma\approx 0.53$ to achieve a comparable privacy guarantee. (See Appendix D). Better analyses of the Gaussian mechanism exist, but do not offer much help. Using the improved analysis in Balle and Wang [2018] to attain the same $\varepsilon$ would require $\sigma\approx 0.34$. Conducting the analysis so that both mechanisms have equivalent privacy loss under zCDP yields $\sigma=0.2$. These are all very large noise magnitudes relative to probabilities in $[0,1]$. [^5] To put independent noise of magnitude $\sigma=0.2$ into perspective: suppose the ground truth next-token prediction is deterministic, i.e., $\bar{\mathbf{p}}=[1,0,...,0]\in\mathbb{R}^{v}$, $v$ = 256128 in the case of Gemma. Now with probability $\geq 0.15$, the noised distribution ${\widetilde{\mathbf{p}}}$ has ${\widetilde{\mathbf{p}}}_{1}<0.8$. Each other $\mathbf{p}_{i}$ is $\geq 0.8$ w.p. $\geq 3\cdot 10^{-5}$ independently. Hence the probability of one of these being promoted to argmax is $\geq 0.15\cdot(1-(1-3\cdot 10^{-5})^{v-1})\approx 0.15$. At this rate, the chance of generating a 30 token span without a corruption is $<1\%$.
Secondly, Tang et al. [2024] generated each token using a different random sample of the sensitive prompts, which is computationally very expensive, as it prevents the use of KV cache-accelerated decoding, since the cache is invalidated upon every resample. While resampling less often would be more practical, Tang et al. [2024] noted that in this case the privacy amplification benefits of subsampling would not be adequately realized, and characterized this limitation as the “main weakness” of their approach. Instead, our method generates each synthetic example using a fixed disjoint subset of the sensitive prompts, allowing us to leverage parallel composition in our analysis, and thus avoid this privacy versus computation tradeoff.
@@ 7 Conclusion
While our work demonstrates that private prediction is a practical technique for privately generating a large volume of high-quality synthetic data, there remains a small gap between our results and the results obtained from privately fine-tuning the parameters of the LLM. Currently, private prediction methods pay a privacy cost for every generated token, while private fine-tuning methods do not. We view correcting this limitation as a very important open problem. Finally, any method for ensuring data privacy will inevitably entail some loss of data utility.
@@
Appendix A Private prediction beats fine-tuning in the limited data regime
We do LoRA fine-tuning with DP-SGD on AGNews1K, with the same setup that beats our method in the full data regime. We sample synthetic data from the fine-tuned model. We also run our private prediction method on AGNews1K. We evaluate performance on 4 and 16 shot in-context learning with GPT-3 babbage-002 (the same experimental setting as §5.1).
@@ Appendix A Private prediction beats fine-tuning in the limited data regime
Table 5: Results on AGNews1K, a 1024-subsample of AGNews. Our method is “pay-as-you-go”, and is capable of generating a few high quality examples for in-context learning in this regime. On the other hand fine-tuning does worse due to the stricter requirement that all future model outputs must be private. We run 16 shot since we only generate 38 examples/since 16 examples fills up the context length for LoRA.
@@ B.1 Logits clipping function
In Figure 3, we compare results for different logits clipping functions. The baseline approach it to clip all logits to the interval $[-c,c]$ before aggregation and softmax – we refer to this as “fixed interval clipping”. Alternatively, we can clip to the range $[\max_{j}\{\mathbf{z}_{j}\}-2c,\max_{j}\{\mathbf{z}_{j}\}]$ and then translate to the interval $[-c,c]$ (Equation 1). In Figure 3 we plot the distortion as a consequence of clipping in terms of L1 error, and find that the latter approach allows us clip more than twice as aggressively, thus improving the privacy guarantee, without compromising utility.

### v2

@@ 4 Privacy analysis
In this section we state formally how Algorithm 1 preserves the privacy of the sensitive prompts it uses to generate synthetic examples.
@@
The proof is in §D and makes use of sharp privacy analyses of: (1) zCDP to approximate DP conversion (Canonne et al., 2020); and (2) zCDP bounds for the exponential mechanism (Cesar and Rogers, 2021).
@@ 5 Experiments
Gemma 1.1 2B IT (Gemma Team, 2024) is the data generator in our main private prediction experiments. We choose it due to its lightweight, open-source JAX implementation [^2] https://github.com/google-deepmind/gemma that makes easy to implement and share sampling algorithms. Tables 1(a) and 1(b) give an overview of datasets and models used.
@@
Gemma 1.1 2B IT Generation; private prediction
@@
(b) Overview of models used in main experiments.
Table 1: Overview of datasets and models used in our main experiments. Datasets are benchmark classification and extraction tasks used in prior work on private synthetic text generation, with the exception of WikiMoviesJSON, which is used for structured data experiments. LaMDA and Gemma are used for synthetic data generation, while the other models are used to measure how useful our synthetic data is for improving accuracy on real test data.
@@
In-context learning (§5.1).
We generate examples to use as in-context exemplars for prompting an LLM. We report downstream accuracy on real test examples, when prompted with synthetic data, on 3 classification tasks (AGNews (Zhang et al., 2015), DBPedia (Zhang et al., 2015), TREC (Voorhees and Tice, 2000)) and 2 extraction tasks (MIT-G, MIT-D (Liu et al., 2012)).
Fine-tuning (§5.2).
We generate synthetic examples to use for fine-tuning a BERT classifier. We report downstream accuracy on real test examples for 3 classification tasks (IMDB (Maas et al., 2011), Yelp (Zhang et al., 2015), AGNews (Zhang et al., 2015)).
Structured data (§5.3).
We generate examples that must adhere to structural constraints to be useful synthetic data. We consider a JSON generation task (WikiMoviesJSON (Rust, 2024)), evaluating structure preservation.
@@ 5.1 In-context learning
Using our method, we generate 90-1500 examples using Gemma 1.1 2B IT. We compare against real examples, and results reported in the prior work of Tang et al. (2024), where they generated 4-shot examples for in-context learning. [^3] It is no longer possible to reproduce their results, due to changes in the OpenAI API since publication: GPT-3 babbage is now deprecated, and it is no longer possible to query for top 100 logprobs, which is required by their method. To evaluate generated synthetic data, we put synthetic examples in the context window before querying with the real test example, as shown in Figure 2.
1Classify the following examples: 2Text: lorem ipsum # synthetic text 1 3Answer: label 4# ... 5Text: sed do eiusm # synthetic text n 6Answer: label 7 8Text: excepteur si # test text 9Answer:
@@
Tang et al. (2024) 4 Tang et al. (2024)* GPT-3 babbage $69.3_{4.8}$ $82.3_{3.7}$ $50.6_{6.9}$ $54.4_{7.0}$
@@
$1$ Tang et al. (2024) 4 Tang et al. (2024)* GPT-3 babbage $64.1_{3.9}$ $81.2_{3.0}$ $50.7_{4.1}$ $46.3_{7.8}$ $69.2_{7.9}$
4 This work Gemma 1.1 2B IT $74.9_{3.8}$ $80.9_{3.6}$ $36.7_{2.2}$ $34.1_{9.3}$ $78.7_{1.9}$
@@
Table 2: In-context learning results with GPT-3 babbage-002. We report mean and standard deviation over 3 random samplings (equally many from each label for classification; fully random for extraction) of synthetic/real data. (*) Note: For the results reported in Tang et al. (2024), they use GPT-3 babbage (now deprecated; we use GPT-3 babbage-002) as the downstream in-context learner, and use the top 100 logprobs for contextual calibration (only top 5 are available now). While not directly comparable, we report their results for context.
@@ Results.
Results are presented in Table 2. Our gains in quantity while maintaining quality are realized in terms of 64-shot in-context learning accuracy. In some cases, we can generate more examples, but we limit ourselves to 64 for these evaluations for cost and efficiency reasons. Our results at 64 shots are comparable to real data at 64 shots. Notably, our synthetic data at 64 shots improves over real data at 4 shots – a rough upper bound on the performance of methods limited to generating 4 examples (e.g., Tang et al. (2024)). We also improve over results reported in Tang et al. (2024) – however as there are differences in the experimental setup, we also report the results of our re-implementation. [^4] Specifically, we use their best hyperparameters (from Appendix E, Table 9 of Tang et al. (2024)) and algorithm, but with our model, prompt, and evaluation setup.
We evaluate with GPT-3 babbage-002 which has a 16K context window. We report results on AGNews, DBPedia, TREC, MIT-G, and MIT-D using the implementation of Zhao et al. (2021). Following the work of Tang et al. (2024), we enable contextual calibration (Zhao et al., 2021) for classification but not extraction tasks. Our evaluation setup is a best-effort reproduction of their setup, which is no longer possible to completely reproduce due to changes to OpenAI API access (see Table 2 caption). Due to cost, we follow prior work (Bertsch et al., 2024; Ratner et al., 2023; Lu et al., 2022; Zhao et al., 2021) and opt to subsample test sets to 250 test examples. We run 3 seeds of sampling of exemplars from synthetic/real data. Additionally, we present a limited set of results on Gemma 2 2/9/27B IT, studying the effect of model size on classification performance in §A.2.
@@ 5.2 Fine-tuning
We use our approach to generate a large quantity of synthetic data for the purposes of fine-tuning 110M BERT-Base models. We consider 3 classification tasks used in prior work on private fine-tuning (Kurakin et al., 2024)), following the exact same evaluation procedure. We omit comparison to prior private prediction work (e.g. (Tang et al., 2024)), as they only generate 4 examples which is insufficient for fine-tuning.
@@ Results.
Main results are presented in Table 3. Across various datasets and privacy levels, we generate between 2.5K (IMDB, $\varepsilon=1$) and 200K (Yelp, $\varepsilon=10$) examples for fine-tuning. Prior work generating fewer than $10$ examples using private prediction were unable to compare with private fine-tuning on these tasks. While there remains a gap between the best fine-tuning and best private inference methods on downstream classification tasks, we achieve reasonable performance, even out-performing or matching the baseline of privately tuning all the parameters in the model reported in Kurakin et al. (2024).
@@
Real data (Kurakin et al., 2024) - $93.7_{0.1}$ $97.6_{0.1}$ $93.7_{0.1}$
Fine-tune (Kurakin et al., 2024) LaMDA 8B $93.2_{0.2}$ $79.1_{1.7}$ $83.9_{0.6}$ $84.0_{0.7}$ $95.9_{0.1}$ $84.1_{0.3}$ $84.6_{0.1}$ $84.2_{0.3}$ $91.1_{0.1}$ $65.7_{2.9}$ $65.3_{2.7}$ $65.1_{5.3}$
@@
Table 3: Results of fine-tuning on real and synthetic data with BERT. We report mean and standard deviation over 3 runs of downstream fine-tuning and evaluation. We compare to results reported in (Kurakin et al., 2024) that fine-tunes a synthetic data generator with DP-SGD. We generate 2.5-200K examples with private prediction, which suffices to train reasonably performing models on.
@@
We additionally consider the limited data regime. In §A.1 we present experiments on AGNews1K, a 1024-subsample of AGNews. Our method, which employs parallel composition, is “pay-as-you-go”, i.e., we can put in a small amount of data to get out a small amount, while preserving quality. On the other hand, fine-tuning based approaches necessarily pay upfront to ensure the model and all future generations are private. This means that without sufficient data, all outputs will be low quality. Results in Table 5 demonstrate that our private prediction method generates more useful examples for in-context learning in this limited data regime.
@@ 5.3 Structured data
For JSON generation, we evaluate on a dataset of information about American movies scraped from Wikipedia (Rust, 2024). Entries contain fields such as title, year, cast, and extract (a short synopsis). We lightly curate the data: we omit uninteresting fields (i.e., thumbnail dimensions and URLs) and remove entries with incomplete entries. We refer to the resulting 34,266 JSON examples with 6 fields as WikiMoviesJSON. We evaluate two criteria: the rate at which output generated constitutes well-formed JSON (Parses (%)), and rate at which the output passes basic schema validation (Validates (%)). This includes checks such as: no extra fields, all required fields are present, values are the correct type, and other custom constraints (e.g. no whitespace in the href field).
@@ 6 Discussion
We believe that our significantly improved performance relative to Tang et al. (2024) is primarily attributable to two algorithmic innovations.
First, for each generated token, Tang et al. (2024) preserve the privacy of the entire distributions from which the token is sampled (by taking argmax), even though only the token itself is included in the synthetic data. By contrast, our method uses a discrete choosing mechanism, the exponential mechanism. As a result, we do not need to maintain a DP version of the entire token distribution to release a single token. This decision leads to significantly lower noise requirements, as a straightforward calculation reveals. Empirically, we obtained good synthetic data quality with $s=250$, $\tau=2$, $c=10$ and $\delta=10^{-6}$. In order to switch to the Gaussian mechanism using its standard $(\varepsilon,\delta)$-DP guarantee, and achieve comparable privacy guarantees we would would require $\sigma\approx 0.53$ to achieve a comparable privacy guarantee. (See §E). Better analyses of the Gaussian mechanism exist, but do not offer much help. Using the improved analysis in Balle and Wang (2018) to attain the same $\varepsilon$ would require $\sigma\approx 0.34$. Conducting the analysis so that both mechanisms have equivalent privacy loss under zCDP yields $\sigma=0.2$. These are all very large noise magnitudes relative to probabilities in $[0,1]$. [^5] To put independent noise of magnitude $\sigma=0.2$ into perspective: suppose the ground truth next-token prediction is deterministic, i.e., $\bar{\mathbf{p}}=[1,0,...,0]\in\mathbb{R}^{v}$, $v$ = 256128 in the case of Gemma. Now with probability $\geq 0.15$, the noised distribution ${\widetilde{\mathbf{p}}}$ has ${\widetilde{\mathbf{p}}}_{1}<0.8$. Each other $\mathbf{p}_{i}$ is $\geq 0.8$ w.p. $\geq 3\cdot 10^{-5}$ independently. Hence the probability of one of these being promoted to argmax is $\geq 0.15\cdot(1-(1-3\cdot 10^{-5})^{v-1})\approx 0.15$. At this rate, the chance of generating a 30 token span without a corruption is $<1\%$.
Secondly, Tang et al. (2024) generated each token using a different random sample of the sensitive prompts, which is computationally very expensive, as it prevents the use of KV cache-accelerated decoding, since the cache is invalidated upon every resample. While resampling less often would be more practical, Tang et al. (2024) noted that in this case the privacy amplification benefits of subsampling would not be adequately realized, and characterized this limitation as the “main weakness” of their approach. Instead, our method generates each synthetic example using a fixed disjoint subset of the sensitive prompts, allowing us to leverage parallel composition in our analysis, and thus avoid this privacy versus computation tradeoff.
@@ 7 Conclusion
While our work demonstrates that private prediction is a practical technique for privately generating a large volume of high-quality synthetic data, there remains a gap between our results and the results obtained from privately fine-tuning the parameters of the LLM. Currently, private prediction methods pay a privacy cost for every generated token, while private fine-tuning methods do not. Finally, any method for ensuring data privacy will inevitably entail some loss of data utility.
@@
Appendix A Additional experiments
A.1 Private prediction beats fine-tuning in the limited data regime
We do LoRA fine-tuning with DP-SGD on AGNews1K, with the same setup that beats our method in the full data regime. We sample synthetic data from the fine-tuned model. We also run our private prediction method on AGNews1K. We evaluate performance on 4 and 16 shot in-context learning with GPT-3 babbage-002 (the same experimental setting as §5.1). Our private prediction approach outperforms the fine-tuning setup that does better in the full data regime.
@@ A.1 Private prediction beats fine-tuning in the limited data regime
Table 5: Results on AGNews1K, a 1024-subsample of AGNews. Our method is “pay-as-you-go”, and is capable of generating a few high quality examples for in-context learning in this regime. On the other hand fine-tuning does worse due to the stricter requirement that all future model outputs must be private. Evaluation setup is the same in §5.1, except here we run 16 instead of 64-shot, because 16 examples produced by the LoRA model fills up the entire context length of babbage-002.
A.2 Effect of model size
We report results on the effect of the data generator’s size on in-context classification performance on DBPedia. Our setup is the same as the experiments in §5.1, with the change that we use the Gemma 2 2/9/27B IT models to get more size variation in the same model family. This necessitated a slight change in the prompt (specifically, we append to the instruction: “No formatting or explanations.” For this evaluation, we see limited improvement due to scale.
$\varepsilon$ Shots Model Acc. (%)
1 4 Gemma 2 2B IT $75.7_{0.8}$
64 $91.2_{0.0}$
4 Gemma 2 9B IT $76.4_{1.2}$
64 $92.4_{1.4}$
4 Gemma 2 27B IT $76.9_{2.2}$
64 $91.9_{1.9}$
Table 6: Results on DBPedia classification. Evaluation setup is the same as in §5.1. We see limited improvement from the increase in model size.
@@ B.1 Logits clipping function
In Figure 3, we compare results for different logits clipping functions. The baseline approach it to clip all logits to the interval $[-c,c]$ before aggregation and softmax – we refer to this as “fixed interval clipping”. Alternatively, we can clip to the range $[\max_{j}\{\mathbf{z}_{j}\}-2c,\max_{j}\{\mathbf{z}_{j}\}]$ and then translate to the interval $[-c,c]$ (Eq. 1). In Figure 3 we plot the distortion as a consequence of clipping in terms of L1 error, and find that the latter approach allows us clip more than twice as aggressively, thus improving the privacy guarantee, without compromising utility.
Appendix C Privacy checklist
We follow privacy reporting guidelines (Ponomareva et al., 2023, Section 5.3.3) to state all information needed to get a complete picture of our privacy guarantee.
  1. 1.
    DP setting. This a central DP guarantee where the service provider is trusted to correctly implement the mechanism.
  2. 2.
    Instantiating the DP definition.
    1. (a)
      Data accesses covered. DP guarantees only apply to a single synthetic dataset generation run. We do not account for hyperparameter tuning.
    2. (b)
      Final mechanism output. The mechanism’s output is the full sequence of sampled tokens of the output synthetic dataset (all synthetic data instances), which can viewed as responses to a sequence of adaptive queries with different input text prefixes.
    3. (c)
      Unit of privacy. Example-level DP, where an example is the full sequence of tokens for one input sensitive example.
    4. (d)
      Adjacency definition for neighboring datasets. Add/remove one example.
  3. 3.
    Privacy accounting details.
    1. (a)
      Type of accounting used. zCDP based accounting.
    2. (b)
      Accounting assumptions. Our analysis assumes independent assignment of each prompt to a batch (Assumption 1). We can fulfill this with hashing (see §4), which corresponds to assigning each prompt to a batch uniformly at random. However, the implementation uses fixed size batches constructed via shuffling.
    3. (c)
      The formal DP statement. We report $(\varepsilon,\delta)$-DP for $\varepsilon=1,3,10$ and $\delta=(\texttt{training\_set\_size})^{-1}$.
  4. 4.
    Transparency and verifiability. We provide a full description of the algorithm and a complete privacy analysis. Our implementation is built off of the official open-source JAX library for Gemma: https://github.com/google-deepmind/gemma. At this time, the implementation is close-sourced.
Discussion.
The full output synthetic dataset of our mechanism indeed satisfies example-level add/remove DP. Below we discuss the more granular aspects of our privacy guarantee that are not captured in the above statement.
First, as a consequence of parallel composition, we note that each individual synthetic example output only incurs privacy cost on the batch used to generate it. Hence, having the mechanism release a subset of the output will only incur privacy cost on a subset of our input dataset. Second, our current privacy guarantee holds with respect to the change of an entire context, which may consist of multiple examples. However, as we do not prescribe a way to group examples into contexts, we cannot a priori identify which examples belong to a single context.